OpenStudy (anonymous):

Evaluate. 6(41 - 21) - 6(20 - 2) =

geerky42 (geerky42):

PEDMAS So evaluate whatever inside parenthesis first.

OpenStudy (anonymous):

6(20) 6(18)

OpenStudy (yanasidlinskiy):

Yes. What's the next step?

OpenStudy (anonymous):

\[6(41 - 21) - 6(20 -2) = 246 - 126 - 120 - 12 = -12\]

OpenStudy (anonymous):

6x20? 6x18?

OpenStudy (anonymous):

Using PEMDAS and distributing the numbers into the parentheses.

OpenStudy (anonymous):

Make sense or should I explain it, @jonxyy ?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

-12 or 12??

OpenStudy (anonymous):

12

geerky42 (geerky42):

@Nibby \(\cdots=246−126−120\color{red}{+12}=12\)